# Exploring Dark Knowledge under Various Teacher Capacities and Addressing Capacity Mismatch

### v1

@@ Abstract
Knowledge Distillation (KD) could transfer the “dark knowledge” of a well-performed yet large neural network to a weaker but lightweight one. From the view of output logits and softened probabilities, this paper goes deeper into the dark knowledge provided by teachers with different capacities. Two fundamental observations are: (1) a larger teacher tends to produce probability vectors that are less distinct between non-ground-truth classes; (2) teachers with different capacities are basically consistent in their cognition of relative class affinity. Abundant experimental studies verify these observations and in-depth empirical explanations are provided. The difference in dark knowledge leads to the peculiar phenomenon named “capacity mismatch” that a more accurate teacher does not necessarily perform as well as a smaller teacher when teaching the same student network. Enlarging the distinctness between non-ground-truth class probabilities for larger teachers could address the capacity mismatch problem. This paper explores multiple simple yet effective ways to achieve this goal and verify their success by comparing them with popular KD methods that solve the capacity mismatch.
Index Terms:
@@ I Introduction
Deploying large-scale neural networks on portable devices with limited computation and storage resources is challenging [1], and efficient architectures such as MobileNets [2, 3] and ShuffleNets [4, 5] have been designed for lightweight deployment. However, the performances of these lightweight networks are usually not comparable to the larger ones. Commonly, second learning [6, 7] or knowledge distillation (KD) [8, 9, 10] could be utilized to transfer the knowledge of a more complex and well-performed network (i.e., the teacher) to the smaller ones (i.e., the student). The dark knowledge in KD is still a mystery that has attracted lots of studies [9, 11, 12, 13], and their goal is to answer the following question: what’s the knowledge that the teacher provides and why they are effective in KD?
@@
In the original KD method [9], the student aims to mimic the teacher’s behavior by minimizing the Kullback-Leibler (KL) divergence between their output probabilities. That is, the logits and softened probabilities, i.e., the inputs to the final softmax operator and the corresponding outputs, are the specific knowledge transferred in KD. With the development of KD methods, the output-level knowledge has been extended to various types [14], including the intermediate features [15, 16, 17, 18, 19, 20], the sample relationships [21, 22, 23, 24, 25, 26, 27], the parameters [28, 29], and the collaborative or online knowledge [10, 30] etc. However, the outputs of neural networks are much easier to visualize, analyze, and understand. Therefore, we focus on the original KD [9] and aim to understand the dark knowledge (i.e., the logits and softened probabilities) provided by the teachers. Unlike previous studies, we majorly study the output-level dark knowledge provided by teachers with various capacities, which receives little attention in previous studies. We first present two significant observations: (1) an over-confident teacher tends to produce probability vectors that are less distinct between non-ground-truth classes; (2) teachers with different capacities are basically consistent in their cognition of relative class affinity. The first observation tells the difference between dark knowledge provided by teachers with different capacities, while the second observation shows the consistency between them.
@@
This paper first explains the reasons for the first observation. Larger teachers generally have powerful feature extractors, making the features of the same class more compact and the features between classes more dispersed. Hence, more complex teachers are over-confident and assign a larger score for the ground-truth class or less varied scores for the non-ground-truth classes. If we use a uniform temperature to scale their logits, the probabilities of non-ground-truth classes are less distinct  [31], further making the distillation process ineffective. This explains the peculiar phenomenon named “capacity mismatch” [32, 33, 34, 35, 36, 31] in KD that a more accurate teacher does not necessarily make better students. Fortunately, the second observation ensures that the dark knowledge from teachers with various capacities is basically consistent in the class relative probabilities. We first provide several definite metrics to verify the observation, including the rank of set overlap, Kendall’s $\tau$, and the Spearman correlation. These metrics are irrelevant to the absolute probability values between classes, which are appropriate for measuring the correlation of relative class affinities. We also present the two observations together via the group classes in CIFAR-100 [37] and a constructed group classification task based on Digits-Five [38].
@@ II-A Dark Knowledge in KD
Quite a few works focus on understanding and revealing the essence and effects of “dark knowledge” in KD, including the empirical studies and theoretical analysis [39, 40, 13, 12, 41, 42, 43, 32, 11, 44]. [11] unifies the dark knowledge with privileged information, and [45] attributes the dark knowledge in vision tasks to the task-relevant and task-irrelevant visual concepts. The dark knowledge in KD is also related to label smoothing (LS) [41, 46, 47]. Some works decompose the dark knowledge of teachers and aim to understand their effectiveness correspondingly. [48] decomposes the dark knowledge into two parts, explaining the teacher’s ground-truth/non-ground-truth outputs as importance weighting and class similarities. [39] decomposes the dark knowledge into three parts including universal knowledge, domain knowledge, and gradient rescaling. The bias-variance decomposition is also utilized to analyze the property of KD [13, 43]. [49] utilizes the standard deviation of the secondary soft probabilities to reflect the quality of the teachers’ dark knowledge. The most related work [31] decomposes the KD into correct guidance, smooth regularization, and class discriminability, which provides novel insights about the key information in the dark knowledge. Few works have examined the relationship of dark knowledge with teacher capacities, and this paper provides detailed observations and analysis about this.
II-B Capacity Mismatch in KD
@@
An intuitive sense after the proposal of KD [9] is that larger teachers could teach students better because their accuracy is higher. However, there exists a peculiar phenomenon named capacity mismatch that excellent teachers can’t completely teach the smaller students well. ESKD [32] first points out the phenomenon and points out that teacher accuracy may poorly predict the student’s performance, i.e., more accurate neural networks don’t necessarily teach better. TAKD [33] solves this problem by introducing an intermediate-sized network (i.e., the teacher assistant) to bridge the knowledge transfer between networks with a large capacity gap. SCKD [34] formulates KD as a multi-task learning problem with several knowledge transfer losses. ResKD [35, 36] utilize the “residual” of knowledge to teach the residual student, and then take the ensemble of the student and residual student for inference. DIST [50] finds that the discrepancy between the student and a stronger teacher may be fairly severe, which disturbs the training process, and they propose a ranking-based loss as the solution. [51] advocates that an intermediate checkpoint will be more appropriate for distillation. [52] explores the strong teachers in few-shot learning. [53] studies the effect of capacity gap on the generated samples in data-free KD. One of our methods is closely related to [31] that proposes Asymmetric Temperature Scaling (ATS) to make the larger networks teach well again. An improved version named Instance-Specific ATS (ISATS) is proposed. Aside from this method, we also propose other simple yet effective solutions from different aspects.
III Background and Preliminaries
@@
We follow the basic notations as introduced in [31]. Specifically, we consider a $C$-class classification problem with $\mathcal{Y}=[C]=\{1,2,\cdots,C\}$. We denote the “logits” as the output scores of a sample $\mathbf{x}$ before applying the softmax function, which is represented as $\mathbf{f}(\mathbf{x})\in\mathbb{R}^{C}$. Correspondingly, the softened probability vector is denoted as $\mathbf{p}(\mathbf{x};\tau)=\text{SF}(\mathbf{f}(\mathbf{x});\tau)$, where $\text{SF}(\cdot;\tau)$ denotes the softmax function with temperature $\tau$. $\mathbf{f}_{y}$ and $\mathbf{p}_{y}$ denote the ground-truth class’s logit and probability, while $\mathbf{g}=[\mathbf{f}_{c}]_{c\neq y}$ and $\mathbf{q}=[\mathbf{p}_{c}]_{c\neq y}$ represent the vector of non-ground-truth classes’ logits and probabilities. $y$ denotes the ground-truth class.
@@
where the upper script “t”/“s” denotes “teacher”/“student” respectively. The dark knowledge in the above distillation equation is the teacher’s label, i.e., $\mathbf{p}^{t}(\tau)$. In this paper, we aim to study the influence of teacher capacity on the teacher’s label. That is, we majorly compare the relationships between $\mathbf{p}^{t_{\text{large}}}$ and $\mathbf{p}^{t_{\text{small}}}$ which are provided by a larger and a smaller teacher, denoted as $t_{\text{large}}$ and $t_{\text{small}}$, respectively.
Commonly, the training and test accuracy of $t_{\text{large}}$ will be higher than that of $t_{\text{small}}$ because the larger teacher has a huge capacity to capture more information. The process of KD has the effect of improving the student network’s performance when compared with its independent training. An explanation for this improvement is that the dark knowledge contained in larger teachers could help the student better capture the semantic information. Hence, to further improve the performance of the student network, a natural idea is replacing the smaller teacher (i.e., $t_{\text{small}}$) with a larger one (i.e., $t_{\text{large}}$). Frustratingly, this leads to a degradation in the performance of the student network. This phenomenon is named the “capacity mismatch” in KD, which is still counter-intuitive, surprising, and unexplored as declared in previous studies [34, 33, 36]. This paper will propose multiple simple yet effective methods to tackle the performance degradation problem.
Figure 1: Visualization of softened probability vectors. The first/second row shows results provided by smaller/larger teachers on the same sample with the same temperature. The first bar shows the ground-truth class while others are non-ground-truth classes. Two observations: (1) the non-ground-truth class probabilities shown in the second row are less varied; (2) the relative class probabilities are nearly the same between the first and second rows.
IV Observations and Explanations
@@
IV-A Observations
An intuitive method to reflect the relationship of dark knowledge produced by teachers with various capacities is visualization. The first step is training teacher networks with different capacities on several benchmarks. Specifically, this paper trains ResNet14 and ResNet110 [54] on the CIFAR-10/CIFAR-100 dataset [37], and trains ResNeXt50-32-4d and ResNeXt101-32-8d [55] on the Stanford Dogs [56] and CUB [57] dataset, and trains WideResNet26-2 and WideResNet50-2 [58] on the TinyImageNet [59] dataset. Then, for each dataset, a random sample is sampled from the first class in the training set and the softened probability vectors produced by different teachers are visualized. Fig. 1 shows the results, where “RES”, “RNX”, and “WRN” are abbreviations for corresponding networks. The utilized temperature is common across the subfigures. Comparing the first and second rows, it is obvious that larger teachers provide less varied probabilities for non-ground-truth classes. Although the variance of non-ground-truth class probabilities differs a lot between teachers, the relative probability values seem to be consistent. For example, the several highest bars between the two rows lie nearly in the same classes. We conclude the two fundamental observations as follows: (1) larger teachers tend to produce probability vectors that are less distinct between non-ground-truth classes; (2) teachers with different capacities are basically consistent in their cognition of relative class affinity. To further support these two observations, we provide explanations and statistical measures in the following.
@@ IV-A Observations
IV-B The Variance of Non-ground-truth Class Probabilities
The explanations for the first observation are provided step by step. First, we train MLP networks on MNIST [60] and train VGG [61] networks on SVHN [62]. These two datasets are relatively simpler classification tasks, and we set the final feature dimensions as 2 and 3 correspondingly for better visualization. We do not use the bias parameters in the last layer. We progressively enlarge the depth and width of MLP networks, and use the total number of parameters to denote the network capacity. The capacity of MLP networks ranges from 0.1M to 0.9M. The VGG capacities are determined by the number of layers, which ranges from 8 to 19. We do not use ReLU [63] activation for the final extracted features. These two datasets have 10 types of digits to identify, and the extracted features are scattered in Fig. 2. With the capacity increases, the features of the same class are more compact, while the class centers between different classes are more dispersed. Notably, the larger networks do not overfit the training dataset, because the test accuracy shown in the figure still becomes higher. This implies that larger DNNs obtain a higher performance but may extract more compact intra-class features and more dispersed inter-class features.
The above demos show low-dimensional features that may not reflect the change in high-dimensional space. Hence, we extend the findings to CIFAR-10 and CIFAR-100, on which we train a series of ResNet, WideResNet, and ResNeXt networks. Because the high-dimensional features are hard to visualize and the dimensional-reduced plots by T-SNE [64] are sensitive to hyper-parameters, we instead calculate several metrics to indirectly verify the change of feature compactness. Specifically, we calculate the feature angle as follows:
@@ IV-B The Variance of Non-ground-truth Class Probabilities
where $\mathbf{h}_{i}$ and $\mathbf{h}_{j}$ are feature vectors of the $i$-th and $j$-th training sample. Then, we define the inter-class and intra-class set as $\{(i,j)\}_{y_{i}=y_{j},\forall i,j}$ and $\{(i,j)\}_{y_{i}\neq y_{j},\forall i,j}$, respectively. The average and standard deviation of angles in the two sets are plotted in Fig. 3. If the teacher becomes more complex, the inter-class feature angle becomes smaller, while the intra-class angle becomes larger. Notably, we use ReLU [63] activation before the classification layer, implying that all of the elements in the extracted features are non-negative. This could already provide some hints for the capacity mismatch phenomenon. As an extreme case, if the features among the same class collapse into a single point, and features among different classes are orthogonal to each other, the inter-class feature angle will be zero and the intra-class one will be 90. Although the training and test accuracy will be $100\%$, the softened probabilities may be one-hot labels, which bring no additional information to the student. An illustration and more detailed discussion can be found in Sect. VII. The inter-class and intra-class distance metrics utilized in linear discriminant analysis [65] or unsupervised discriminant projection [66] show the same trend. We omit the display of the results based on these distance metrics because the trends of feature angles are vivid enough to show how the feature compactness changes.
Figure 3: The change of inter-class and intra-class feature angle under different teacher capacities. The x-axis in each subfigure shows the teacher capacity.
@@
According to the softmax properties in classification problems, the classification weight of each class is pulled towards the features from the same class while pushed away from the features of other classes [67]. That is, the change tendency of features could partially reflect the tendency of classification weights in the final layer. An ideal case is that the learned classification weights converge to the feature centers. If the intra-class feature angle becomes larger, the feature of a specific sample will be far away from the classification weights of the other classes. Similarly, the feature of a specific sample will be closer to the corresponding class’s classification weight. This will lead to a larger logit for the ground-truth class. If the intra-class features are nearly orthogonal, the logits for non-ground-truth classes will be less varied. The illustration can be found in Fig. 4, where the larger teacher extracts more compact inter-class features and more dispersed intra-class features. The triangle represents a sample feature in the $1$-st class, and the larger teacher tends to give a larger target logit (i.e., $\mathbf{f}_{1}=\mathbf{h}^{T}\mathbf{w}_{1}$) and less varied non-ground-truth class logits (i.e., $\mathbf{g}=[\mathbf{h}^{T}\mathbf{w}_{c}]_{c\in\{2,3,4,5\}}$). A larger target logit or less varied non-ground-truth class logits will both lead to less varied non-ground-truth class probabilities after applying softmax function [31]. The step-by-step analysis and visualization provide explanations for the first observation that larger teachers tend to produce probability vectors that are less varied between non-ground-truth classes.
IV-C The Relative Class Affinities

### v2

@@ Abstract
Knowledge Distillation (KD) could transfer the “dark knowledge" of a well-performed yet large neural network to a weaker but lightweight one. From the view of output logits and softened probabilities, this paper goes deeper into the dark knowledge provided by teachers with different capacities. Two fundamental observations are: (1) a larger teacher tends to produce probability vectors with lower distinction among non-ground-truth classes; (2) teachers with different capacities are basically consistent in their cognition of relative class affinity. Through abundant experimental studies we verify these observations and provide in-depth empirical explanations to them. We argue that the distinctness among incorrect classes embodies the essence of dark knowledge. A larger and more accurate teacher lacks this distinctness, which hampers its teaching ability compared to a smaller teacher, ultimately leading to the peculiar phenomenon named "capacity mismatch". Building on this insight, this paper explores multiple simple yet effective ways to address capacity mismatch, achieving superior experimental results compared to previous approaches.
doi:
articletype: RESEARCH ARTICLE
[…]
@@ 1 Introduction
Deploying large-scale neural networks on portable devices with limited computation and storage resources is challenging [1], and efficient architectures such as MobileNets [2, 3] and ShuffleNets [4, 5] have been designed for lightweight deployment. However, the performances of these lightweight networks are usually not comparable to the larger ones. Commonly, second learning [6, 7] or knowledge distillation (KD) [8, 9, 10] could be utilized to transfer the knowledge of a more complex and well-performed network (i.e., the teacher) to the smaller ones (i.e., the student). Both are manifestations of pre-learning transfer, with KD focusing more on extracting the dark knowledge from the teacher model. However, the dark knowledge in KD is still a mystery that has attracted lots of studies [9, 11, 12, 13], and their goal is to answer the following question: what’s the knowledge that the teacher provides and why they are effective in KD?
@@
In the original KD method [9], the student aims to mimic the teacher’s behavior by minimizing the Kullback-Leibler (KL) divergence between their output probabilities. That is, the logits and softened probabilities, i.e., the inputs to the final softmax operator and the corresponding outputs, are the specific knowledge transferred in KD. With the development of KD methods, the output-level knowledge has been extended to various types [14], including the intermediate features [15, 16, 17, 18, 19, 20], the sample relationships [21, 22, 23, 24, 25, 26, 27], the parameters [28, 29], and the collaborative or online knowledge [10, 30] etc. Among these types, the output logits of neural networks are much easier to visualize, analyze, and understand. Therefore, we focus on the original KD [9] and aim to understand the dark knowledge (i.e., the logits and softened probabilities) provided by the teachers. Unlike previous studies, we majorly study the output-level dark knowledge provided by teachers with various capacities, which receives little attention in previous studies. We first present two significant observations: (1) an over-confident teacher tends to produce probability vectors that are less distinct between non-ground-truth classes; (2) teachers with different capacities are basically consistent in their cognition of relative class affinity. The first observation tells the difference between dark knowledge provided by teachers with different capacities, while the second observation shows the consistency between them.
@@
This paper first explains the reasons for the first observation. Larger teachers generally have powerful feature extractors, making the features of the same class more compact and the features between classes more dispersed. Hence, more complex teachers are over-confident and assign a larger score for the ground-truth class or less varied scores for the non-ground-truth classes. If we use a uniform temperature to scale their logits, the probabilities of non-ground-truth classes are less distinct  [31], further making the distillation process ineffective. This explains the peculiar phenomenon named “capacity mismatch" [32, 33, 34, 35, 36, 31] in KD that a larger and more accurate teacher does not necessarily make better students. Fortunately, the second observation ensures that the dark knowledge from teachers with various capacities is basically consistent in the class relative probabilities. We first provide several definite metrics to verify the observation, including the rank of set overlap, Kendall’s $\tau$, and the Spearman correlation. These metrics are irrelevant to the absolute probability values between classes, which are appropriate for measuring the correlation of relative class affinities. We also present the two observations together via the group classes in CIFAR-100 [37] and a constructed group classification task based on Digits-Five [38].
@@ 2.1 Dark Knowledge in KD
Quite a few works focus on understanding and revealing the essence and effects of “dark knowledge" in KD, including the empirical studies and theoretical analysis [39, 40, 13, 12, 41, 42, 43, 32, 11, 44]. [11] unifies the dark knowledge with privileged information, and [45] attributes the dark knowledge in vision tasks to the task-relevant and task-irrelevant visual concepts. The dark knowledge in KD is also related to label smoothing (LS) [41, 46, 47, 48]. Some works decompose the dark knowledge of teachers and aim to understand their effectiveness correspondingly. [49] decomposes the dark knowledge into two parts, explaining the teacher’s ground-truth/non-ground-truth outputs as importance weighting and class similarities. [39] decomposes the dark knowledge into three parts including universal knowledge, domain knowledge, and gradient rescaling. The bias-variance decomposition is also utilized to analyze the property of KD [13, 43]. [50] utilizes the standard deviation of the secondary soft probabilities to reflect the quality of the teachers’ dark knowledge. Recently, [51] proposed an innovative perspective, suggesting that the knowledge of the teacher stems from containing less noise. The most related work [31] decomposes the KD into correct guidance, smooth regularization, and class discriminability, which provides novel insights about the key information in the dark knowledge. Few works have examined the relationship of dark knowledge with teacher capacities, and this paper provides detailed observations and analysis about this.
2.2 Capacity Mismatch in KD
@@
An intuitive sense after the proposal of KD [9] is that larger teachers could teach students better because their accuracy is higher. However, there exists a peculiar phenomenon named capacity mismatch that excellent teachers can’t completely teach the smaller students well. ESKD [32] first points out the phenomenon and points out that teacher accuracy may poorly predict the student’s performance, i.e., more accurate neural networks don’t necessarily teach better. However, no research so far has provided an accurate explanation or even a quantitative definition for this phenomenon. Previous works have only been made to propose solutions from various perspectives. TAKD [33] solves this problem by introducing an intermediate-sized network (i.e., the teacher assistant) to bridge the knowledge transfer between networks with a large capacity gap. SCKD [34] formulates KD as a multi-task learning problem with several knowledge transfer losses. ResKD [35, 36] utilize the “residual" of knowledge to teach the residual student, and then take the ensemble of the student and residual student for inference. DIST [52] finds that the discrepancy between the student and a stronger teacher may be fairly severe, which disturbs the training process, and they propose a ranking-based loss as the solution. [53] advocates that an intermediate checkpoint will be more appropriate for distillation. [54] explores the strong teachers in few-shot learning. [55] studies the effect of capacity gap on the generated samples in data-free KD. Recently, [56] demonstrated that sharing the same temperature between the teacher and student leads to a logit shift, which is a side-effect in KD. One of our methods is closely related to [31] that proposes Asymmetric Temperature Scaling (ATS) to make the larger networks teach well again. An improved version named Instance-Specific ATS (ISATS) is proposed. Aside from this method, we also propose other simple yet effective solutions from different aspects.
3 Background and Preliminaries
@@
We follow the basic notations as introduced in [31]. Specifically, we consider a $C$-class classification problem with $\mathcal{Y}=[C]=\{1,2,\cdots,C\}$. We denote the “logits" as the output scores of a sample $\mathbf{x}$ before applying the softmax function, which is represented as $\mathbf{f}(\mathbf{x})\in\mathbb{R}^{C}$. Correspondingly, the softened probability vector is denoted as $\mathbf{p}(\mathbf{x};\tau)=\text{SF}(\mathbf{f}(\mathbf{x});\tau)$:
$\text{SF}(\mathbf{f}(\mathbf{x});\tau)=\exp\left(\mathbf{f}(\mathbf{x})/\tau\right)/Z(\tau),$ (1)
where $Z(\tau)=\sum_{i=1}^{C}\exp(f_{i}(\mathbf{x})/\tau)$. $f_{i}(\mathbf{x})$ is a scalar which refers to the logit of the $i$-th class in $\mathbf{f}(\mathbf{x})$. $\mathbf{f}_{y}$ and $\mathbf{p}_{y}$ denote the ground-truth class’s logit and probability, while $\mathbf{g}=[\mathbf{f}_{c}]_{c\neq y}$ and $\mathbf{q}=[\mathbf{p}_{c}]_{c\neq y}$ represent the vector of non-ground-truth classes’ logits and probabilities. $y$ denotes the ground-truth class.
@@
where the upper script “t”/“s” denotes “teacher”/“student” respectively. We utilize $\lambda\in[0,1]$ as a hyperparameter to balance the weights between the two training stages. The dark knowledge in the above distillation equation is the teacher’s label, i.e., $\mathbf{p}^{t}(\tau)$. In this paper, we aim to study the influence of teacher capacity on the teacher’s label. That is, we majorly compare the relationships between $\mathbf{p}^{t_{\text{large}}}$ and $\mathbf{p}^{t_{\text{small}}}$ which are provided by a larger and a smaller teacher, denoted as $t_{\text{large}}$ and $t_{\text{small}}$, respectively.
The process of KD has the effect of improving the student network’s performance when compared with its independent training. An explanation for this improvement is that the dark knowledge contained in larger teachers could help the student better capture the semantic information. Commonly, the training and test accuracy of $t_{\text{large}}$ will be higher than that of $t_{\text{small}}$ because the larger teacher has a huge capacity to capture more information. Hence, to further improve the performance of the student network, a natural idea is replacing the smaller teacher (i.e., $t_{\text{small}}$) with a larger one (i.e., $t_{\text{large}}$). Frustratingly, this leads to a degradation in the performance of the student network. This phenomenon is named the “capacity mismatch” in KD, which is still counter-intuitive, surprising, and unexplored as declared in previous studies [34, 33, 36]. This paper will propose multiple simple yet effective methods to tackle the performance degradation problem.
Figure 1: Visualization of softened probability vectors. (Up) Results provided by smaller teacher on various datasets; (Bottom) Results provided by larger teacher on various datasets. Each column’s two rows represent the same sample with the same temperature. The first bar shows the ground-truth class while others are non-ground-truth classes. Two observations: (1) the non-ground-truth class probabilities shown in the second row are less varied; (2) the relative class probabilities are nearly the same between the first and second rows.
4 Observations and Explanations
@@
4.1 Observations
An intuitive method to reflect the relationship of dark knowledge produced by teachers with various capacities is visualization. The first step is training teacher networks with different capacities on several benchmarks. Specifically, we train ResNet14 and ResNet110 [57] on the CIFAR-10/CIFAR-100 dataset [37], ResNeXt50-32-4d and ResNeXt101-32-8d [58] on the Stanford Dogs [59] and CUB [60] dataset, and WideResNet26-2 and WideResNet50-2 [61] on the TinyImageNet [62] dataset. We maximize the size difference between the two models on each dataset to make the experimental effects more pronounced. Then, for each dataset, a random sample is sampled from the first class in the training set and the softened probability vectors produced by different teachers are visualized. Fig. 1 shows the results, where “RES", “RNX", and “WRN" are abbreviations for corresponding networks. The utilized temperature is common across the subfigures. Comparing the first and second rows, it is obvious that larger teachers provide less varied probabilities for non-ground-truth classes. Although the variance of non-ground-truth class probabilities differs a lot between teachers, the relative probability values seem to be consistent. For example, the several top-$K$ bars between the two rows lie nearly in the same classes. We conclude the two fundamental observations as follows: (1) larger teachers tend to produce probability vectors that are less distinct among non-ground-truth classes; (2) teachers with different capacities are basically consistent in their cognition of relative class affinity. To further support these two observations, we provide explanations and statistical measures in the following.
@@ 4.1 Observations
4.2 The Variance of Non-ground-truth Class Probabilities
The explanations for the first observation are provided step by step.
Low-dimensional features. First, we train MLP networks on MNIST [63] and train VGG [64] networks on SVHN [65]. These two datasets are relatively simpler classification tasks, and we set the final feature dimensions as 2 and 3 correspondingly for better visualization. We do not use the bias parameters in the last layer. We progressively enlarge the depth and width of MLP networks, and use the total number of parameters to denote the network capacity. The capacity of MLP networks ranges from 0.1M to 0.9M. The VGG capacities are determined by the number of layers, which ranges from 8 to 19. We do not use ReLU [66] activation for the final extracted features. These two datasets have 10 types of digits to identify, and the extracted features are scattered in Fig. 2. With the capacity increases, the features of the same class are more compact, while the class centers between different classes are more dispersed. Notably, the larger networks do not overfit the training dataset, because the test accuracy shown in the figure still becomes higher. This implies that larger DNNs obtain a higher performance but may extract more compact intra-class features and more dispersed inter-class features.
High-dimensional features. The above demos show low-dimensional features that may not reflect the change in high-dimensional space. Hence, we extend the findings to CIFAR-10 and CIFAR-100, on which we train a series of ResNet, WideResNet, and ResNeXt networks. Because the high-dimensional features are hard to visualize and the dimensional-reduced plots by T-SNE [67] are sensitive to hyper-parameters, we instead calculate several metrics to indirectly verify the change of feature compactness. Specifically, we calculate the feature angle as follows:
@@ 4.2 The Variance of Non-ground-truth Class Probabilities
where $\mathbf{h}_{i}$ and $\mathbf{h}_{j}$ are feature vectors of the $i$-th and $j$-th training sample. Then, we define the inter-class and intra-class set as $\{(i,j)\}_{y_{i}=y_{j},\forall i,j}$ and $\{(i,j)\}_{y_{i}\neq y_{j},\forall i,j}$, respectively. The average and standard deviation of angles in the two sets are plotted in Fig. 3. If the teacher becomes more complex, the inter-class feature angle becomes smaller, while the intra-class angle becomes larger. Notably, we use ReLU [66] activation before the classification layer, implying that all of the elements in the extracted features are non-negative. This could already provide some hints for the capacity mismatch phenomenon. As an extreme case, if the features among the same class collapse into a single point, and features among different classes are orthogonal to each other, the inter-class feature angle will be zero and the intra-class one will be 90. Although the training and test accuracy will be $100\%$, the softened probabilities may be one-hot labels, which bring no additional information to the student. An illustration and more detailed discussion can be found in Sect. 7. The inter-class and intra-class distance metrics utilized in linear discriminant analysis [68] or unsupervised discriminant projection [69] show the same trend. We omit the display of the results based on these distance metrics because the trends of feature angles are vivid enough to show how the feature compactness changes.
Figure 3: The change of inter-class and intra-class feature angle under different teachers. (a) ResNet teachers; (b) WideResNet teachers; (c) ResNeXt teachers
@@
Visualizaion of classes. According to the softmax properties in classification problems, the classification weight of each class is pulled towards the features from the same class while pushed away from the features of other classes [70]. That is, the change tendency of features could partially reflect the tendency of classification weights in the final layer. An ideal case is that the learned classification weights converge to the feature centers. If the intra-class feature angle becomes larger, the feature of a specific sample will be far away from the classification weights of the other classes. Similarly, the feature of a specific sample will be closer to the corresponding class’s classification weight. This will lead to a larger logit for the ground-truth class. If the intra-class features are nearly orthogonal, the logits for non-ground-truth classes will be less varied. The illustration can be found in Fig. 4, where the larger teacher extracts more compact inter-class features and more dispersed intra-class features. The triangle represents a sample feature in the $1$-st class, and the larger teacher tends to give a larger target logit (i.e., $\mathbf{f}_{1}=\mathbf{h}^{T}\mathbf{w}_{1}$) and less varied non-ground-truth class logits (i.e., $\mathbf{g}=[\mathbf{h}^{T}\mathbf{w}_{c}]_{c\in\{2,3,4,5\}}$). A larger target logit or less varied non-ground-truth class logits will both lead to less varied non-ground-truth class probabilities after applying softmax function [31].
The step-by-step analysis and visualization provide explanations for the first observation that larger teachers tend to produce probability vectors that are less varied between non-ground-truth classes.
4.3 The Relative Class Affinities